OpenStudy (anonymous):

Topic: \(Calculus \ 2\), polar conversions Q: Identify name & Cartesian equation for: r = 2\(\tan \theta \sec \theta\) Any graphing calculator can quickly show this is a parabola the one I've sketched below, but I'm trying to understand the process here. These are the facts I know of: \(r=\sqrt{x^2+y^2}\) \(\theta = \tan^{-1}(\frac{y}{x})\) x = r cos θ y = r sin θ What's the trick here? Can somebody show me? (I'd be more than happy to give out a medal if you can do so) There are a few other problems related to this one I'm working on so a technique to be learned is the goa

OpenStudy (anonymous):

\[\large \sqrt{x^2+y^2}=\tan(\tan^{-1}\frac{y}{x})\sec(\tan^{-1}\frac{y}{x})\] But then um... lol what?